OpenStudy (anonymous):

I can't seem to get the correct answer for solving this derivative using the product rule...

OpenStudy (anonymous):

\[h (t) = \sqrt{t} (1 - t ^{2})\]

OpenStudy (anonymous):

\[h \prime (t) = (\sqrt{t}) (-2t) + ( 1 - t ^{2}) (1/2 t ^{-1/2})\] This is what I got after plugging in the product rule

myininaya (myininaya):

looks good

OpenStudy (anonymous):

Arey .... \[h(t) = \sqrt{t} - t^{5/2} \]\[\frac{d h(t)}{dt} = \frac{1}{2} \times t^{1/2 -1 } + \frac{5}{2} \times t ^{\frac{5}{2} - 1} \]\[h'(t) = \frac{1}{2 \times \sqrt{t}} + \frac{5 t^{3/2}}{2}\]

OpenStudy (anonymous):

Where did the \[-t ^{5/2}\] come from?

OpenStudy (anonymous):

t^1/2 x t^ 2 = t^1/2+2 = t^5/2

OpenStudy (anonymous):

Oh okay, I didn't know you could combine \[\sqrt{t} , and -t ^{2}\] like that. I'll have to keep that in mind. Thanks

myininaya (myininaya):

remember a(b+c)=ab+ac

myininaya (myininaya):

but what you had done above using product rule was fine too

OpenStudy (anonymous):

myininaya do you know how can we drive derivative of inverse functions

OpenStudy (anonymous):

should i post it as a question there you can tell me

myininaya (myininaya):

yes i think i remember how ok ishaan give me a few i would rather write on paper and scan and post

OpenStudy (anonymous):

thanks

myininaya (myininaya):

wrong attachment

OpenStudy (anonymous):

thanks ...I tried it with limits and all... got confused ...thank you very much : )

myininaya (myininaya):

the trick is too draw a picture well some peeps might not need a picture but i certainly do

myininaya (myininaya):

i know you probably get it by now so i will stop lol its kindof fun though

OpenStudy (anonymous):

yeah i got it thanks again it's a good solution : )

myininaya (myininaya):

i derived some of inverse trig things above

myininaya (myininaya):

let me know if you have any questions,chaise

OpenStudy (chaise):

Thanks a bunch! I'm afraid I'm not aware of what csc and sec are. Should I know what these are? I don't want to burden by asking, cause I understand how difficult it can be to explain, but I will try and derive as much knowledge from your examples and also ask my maths teacher to further explain.

myininaya (myininaya):

the trick is to make a right traingle and say we are wanting to find the derivative of y=arcsin(x) then taking sin of both sides we have siny=x so y is an angle and we have siny=x/1 so that means the opposite side of that angle has measure x and the hyp. has measure 1 so we can find the other side by using Pythagorean thm sqrt{1^2-x^2}=sqrt{1-x^2} so we have siny=x we can use implicit differentiation to find y' cosy*y'=1 y'=1/cosy but what is this in terms of x well we can referred to our triangle right? cosy means adj/hyp hyp=1 and we found the adjacent side earlier using the Pythagorean thm

myininaya (myininaya):

sec means hyp/adj csc means hyp/opp

myininaya (myininaya):

cos means adj/hyp so you can say sec means 1/cos

OpenStudy (chaise):

These new terms are so foreign to me. Why have I never be explained these before? Is there one your missing? Like.. a tan function one?

myininaya (myininaya):

yes you can also do the above process to find the derivative of tan inverse and csc inverse

OpenStudy (chaise):

cool. thank you so much!

myininaya (myininaya):

if we have y=arctan(x), then tany=x right? tany means opposite/adjacent right? we have tany=x/1 so what side is missing and what is its measure?

OpenStudy (chaise):

we are missing the opposite side and its measure is 1/tany? Is this right?

myininaya (myininaya):

no ok tany means opposite/adjacent right?

myininaya (myininaya):

so we are missing the hypotenuse

OpenStudy (chaise):

tan(y) = x/1 tan y = opp/adj? The opposite side is x and the adj is 1?

OpenStudy (chaise):

Oooh. we are missing hyptonusel but the x is unknown. I see now. sorry.

myininaya (myininaya):

| \ | \ h x| y\ --- 1 right? so we need to know what h is we can find measure h by using pythagorean thm h=sqrt{x^2+1}

OpenStudy (chaise):

yes. im following.

myininaya (myininaya):

ok so anyways we will come back to the triangle we have tany=x right? so taking derivative of both sides we get y' sec^2(y)=1

myininaya (myininaya):

to find y' which is our objective we divide both sides by sec^2(y) y'=1/(sec^2y)

myininaya (myininaya):

but we need the derivative in terms of x

myininaya (myininaya):

using our traingle what is secy in terms of x

OpenStudy (chaise):

Aah. and then you have to change the derivative back to terms of x, and yes I see now. :) THANK YOU SO MUCH.